OpenStudy (anonymous):

You are a manufacturer of a certain type of watch, and you’ve calculated that it costs $600 to make 35 watches, and it costs $715 to make 50 watches. Assuming that the relationship between cost and the number of watches you make (volume) is linear, write the equation that relates cost to the number of watches you make and use the equation to calculate how much it would cost to produce 42 watches. Round your answer to the nearest dollars and cents. Be sure to use the dollar sign and the decimal point for the correct answer

OpenStudy (anonymous):

since it is linear we gonna have to evaluate on both cases. case one: $600=35 watched and x= 42watches we find x=600*42/35=720 dollar case two: $715=50 watches x= 42 watches we find x=715*42/50 =$715

OpenStudy (anonymous):

so the first one will need $720 and next one will need $715